A rock is thrown horizontally from the top of a building with an initial speed of v = 10.1 m/s. If it lands d = 57.1 m from the base of the building, how high is the building?


1
Expert's answer
2021-10-06T15:52:13-0400

The speed in horizontal direction remains unchanged during the falling. Thus, the time of falling is given as follows:


"t = \dfrac{d}{v_h}"

where "d = 57.1m" is the range, and "v_h = 10.1m\/s" is the horizontal speed.

The distance travelled in vertical direction (height of the building) is given by the following kinematic equation:


"h = \dfrac{gt^2}{2}"

where "g = 9.81m\/s^2" is the gravitational acceleration. Substituting the expression for "t", obtain:


"h = \dfrac{gd^2}{2v_h^2} = \dfrac{9.81\cdot 57.1^2}{2\cdot 10.1^2} \approx 157m"

Answer. 157 m.